# Robot Localization aided by Quantum Algorithms

### v1

@@ 1 Introduction
Localization is a vital aspect of mobile robotics, enabling robots to navigate their environment efficiently and avoid obstacles. Without localization, mobile robots would be unable to determine their position and orientation, making it challenging to plan a path or make informed decisions about their movement (Olson (2000)). Localization allows mobile robots to create an internal map of their environment, which is essential for tasks such as surveying, manipulation, inspection, and delivery (Huang and Lin (2023)). In fact, localization is what enables mobile robots to perform tasks autonomously, making informed decisions about their actions and movements without human intervention.
@@
The quality of localization is heavily dependent on the generation of accurate maps, which is a computationally intensive task. Probabilistic localization methods, such as the Adaptive-Monte Carlo localization (AMCL) algorithm, have been widely used in mobile robotics due to their accuracy and robustness (Kristensen and Jensfelt (2003)). However, these methods can be computationally demanding, especially when dealing with large maps or high-resolution sensor data. AMCL, in particular, uses a combination of sensor data and prior map knowledge to determine the probable location of a robot on a given map, but its computation complexity is proportional to the area of the grid of the map (Alshikh Khalil and Hatem (2022)).
@@
Recently, the integration of light detection and ranging (LiDAR) sensors has improved the accuracy of localization methods, but the computational requirements remain a challenge (Huang and Lin (2023)). To overcome this challenge, researchers have started exploring the application of quantum computing in robotics, which has the potential to revolutionize the field by providing faster and more efficient computation (Marek (2020)).
@@
Quantum computing is a new paradigm for computing that has shown significant promise in various fields, including cryptography, optimization, and machine learning. Its application in robotics is still in its early stages, but it has the potential to provide breakthroughs in areas such as localization, planning, and control (Petschnigg et al. (2019)). Quantum-inspired algorithms, such as Grover’s search, have been shown to provide computational speedups in certain tasks, making them an attractive solution for robotic applications (Tandon et al. (2017)). Different proposals have been subject to research, such as planning (Chella et al. (2023)), but new approaches that benefit from the quantum advantage have not yet fully explored.
@@
This paper aims to evaluate the introduction of quantum computing in robotics, focusing on the application of quantum algorithms in robot localization. Specifically, it explores the use of Grover’s search algorithm in AMCL to improve the efficiency of localization in mobile robots. While previous studies have explored the application of quantum planners in robotic navigation, their primary focus has been on utilizing quantum algorithms to enhance decision-making optimization rather than integrating quantum computing into the localization module itself (Chella et al. (2022)). Our approach aims to bridge this gap by introducing a pragmatic approach to robot localization that can be effectively implemented in real-world scenarios.
@@ 1.1 Research Objectives
Quantum computing has the potential to bring significant benefits to robotics science by enhancing computational power, optimizing algorithms, and enabling new types of processing that are not possible with classical computers (Montanaro (2015)). Quantum algorithms leverage the principles of quantum mechanics to perform certain computations more efficiently than classical algorithms.
Grover’s algorithm is one of the most famous quantum algorithms introduced by Lov Grover in 1996 (Grover (1996)). It is a quantum search algorithm that finds an element in an unsorted database of N items in O($\sqrt{N}$) time, compared to classical algorithms that require O(N) time. Grover’s algorithm is suitable for applications involving unsorted database search, cryptography, and combinatorial optimization, where it can provide a quadratic speedup over classical algorithms (Guo (2023)).
@@ 2 Background
Effective navigation is a fundamental capability for mobile robots, relying on the successful integration of four essential components: sensor-based perception, self-localization, decision-making, and precise motion control. Among these components, self-localization is crucial, as it enables the robot to determine its position and orientation, providing a comprehensive understanding of its spatial disposition within its environment.
@@
The widespread use of Global Positioning Systems (GPS) for outdoor navigation has become a cornerstone of modern technology, enabling devices to determine their location with relative ease. However, GPS signals are not immune to errors and uncertainties, particularly in certain environments. Atmospheric interference, multipath effects, and satellite geometry can all compromise the accuracy of GPS signals, resulting in positioning errors of up to 10-20 meters. Although this level of uncertainty may be tolerable for some applications, it is unacceptable for many robotic systems that require precise navigation to function effectively.
The limitations of GPS become even more pronounced in environments where signals are unavailable or unreliable. Indoor spaces, urban canyons, and areas with heavy vegetation can all disrupt or block GPS signals, hindering a robot’s ability to maintain a consistent and accurate estimate of its position. This can lead to reduced navigation performance, increased risk of collisions, and compromised overall system reliability.
In opposition to such global localization systems, sensor-based localization methods have emerged as a reliable and accurate means of navigation for mobile robots. By integrating data from various sensors, such as lidar, cameras, and inertial measurement units, robots can build a more comprehensive and accurate understanding of their surroundings. This multisensory approach enables robots to perceive their environment in greater detail, allowing for more effective navigation and control.
The use of multiple sensors provides a more nuanced understanding of the environment, including the location of obstacles, terrain features, and other relevant factors. This information can be used to improve navigation accuracy, reduce the risk of collisions, and improve overall system reliability. Sensor-based localization methods can be designed to function in a wide range of environments, including those where GPS signals are unavailable or unreliable. Despite the advantages of sensor-based localization methods, there is still room for improvement. The integration of data from multiple sensors can be a complex task, requiring sophisticated algorithms and processing capabilities. In addition, the accuracy and reliability of sensor-based localization methods can be affected by various factors, including sensor noise, calibration errors, and environmental conditions. Therefore, ongoing research is focused on developing more advanced sensor-based localization methods that can provide even greater accuracy, reliability, and robustness in a wide range of environments.
In these sensor-based localization methods for mobile robots, there are two main complementary techniques (Qu et al. (2018)): landmark-based absolute localization systems and probabilistic localization methods.
Landmark-based robot localization systems enable robots to determine their location within an environment by identifying and tracking distinctive features or landmarks. These landmarks can be natural or artificial and are typically extracted from sensor data, such as images, lidar scans, or GPS readings.
On the other hand, probabilistic localization methods, such as Bayesian filtering and Monte Carlo localization (Zhang and Zapata (2009)), maintain a probability distribution over possible locations and update this distribution based on sensor data and motion models. These methods can effectively handle noisy sensor data, ambiguous landmarks, and dynamic environments, providing a robust and accurate estimate of the robot’s location.
However there are two aspects where all localization techniques agree:
  • All of them need some map (which means that all of them require some way of encoding the map’s information)
  • All of them require some way to search the robot pose in such map
It is worth mentioning that depending on the map’s size, both the encoding and the search process can be quite complex.
Although these classical computing-based localization methods have shown significant progress, the increasing complexity of robotic applications demands more efficient and robust techniques. The emergence of quantum computing offers new opportunities to enhance localization methods, enabling faster processing and more accurate positioning. In the near future, quantum computing is expected to revolutionize the field of robotics in general, and localization in particular, enabling robots to navigate and interact with their environment.
@@ 2.1 Map encoding
Map encoding (encoding the map information into a form that can be manipulated by the robot’s CPU) is a required step in any localization technique. In the last decades, robotic research has explored and developed different mapping technologies (Thrun (2003)), (Lluvia et al. (2021)), being the "occupancy grid" approach one of the dominant paradigms in the robotics field (Collins et al. (2007)).
The core concept of the occupancy grid is to model the environment as a regularly spaced grid, where each cell is a binary random variable indicating whether or not an obstacle exists at that specific location.
A possible way to encode the map could be to assign different states to the cells. For example: "1" for cells that contain an obstacle, and "0" for cells without obstacles (cells than can be traversed by a robot)
Using this encoding, we can describe the map as a matrix. Let’s take for example a simple map as shown in the next Figure 1:
Figure 1: : Simple 4x4 grid that represents a map with 5 obstacles (black boxes)
In the case of the previous map defined in Figure 1, the matrix would be as:
$\begin{bmatrix}1&0&1&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&0&0&1\\ \end{bmatrix}$ (1)
One limitation of this approach is that the information is limited only to where the obstacles are located, without any other extra data. Historically, occupancy grids have been extended to include more information, hence creating a "costmap". The costmap is a regular 2D grid of the environment, in which information from a series of sensor processing plugins, also called costmap layers, is stored (Ferguson and Likhachev (2008)), (Han et al. (2017)).
In the scope of this paper, we propose to use a costmap to encode the environment of the robot. In robotics, a costmap is a grid-based representation of the environment used to aid navigation, path planning, and obstacle avoidance. Each cell in the costmap is assigned a numerical value or "cost" that represents the difficulty or risk associated with traveling through that specific area. This contrasts with occupancy grids, which are typically binary and indicate whether a location is occupied or free. The costmap makes it possible to encode more information than just the obstacles. In this costmap based encoding obstacles, free spaces, and areas of varying risk can be represented.
The values in a costmap are determined by several factors, such as the presence of obstacles, proximity to other obstacles, and other environmental conditions. Costmaps can be created using different algorithms, but in almost all of them is common to represent obstacles inflated by the inscribed radius of the robot, to minimize collision risk (Wen and Yang (2021)).
In this paper, we explore this concept, proposing a simple algorithm that creates a costmap that tries not only to identify the specific location of obstacles, but also the presence of nearby obstacles.
The proposed algorithm for the cost map has three steps: a first step where for every obstacle its influence in nearby cell is computed is presented, a second step where for each cell, a sum of such influences is calculated, and a final third step where the cost values are normalized to compact the information (use less bits in its representation). These steps are presented in Algorithm 1:
Algorithm 1 Simple costmap algorithm
$B_{1}\dots B_{N}$ as a list of N obstacles
$C_{1}\dots C_{M}$ a a list of M cells in a $n\times n$ grid, being $M=n\times n$
@@
$Costmap~{}value~{}of~{}cell~{}C_{1}\dots C_{M}\leftarrow 0$
for $k\leftarrow 1$ to $N$ do
@@
         $Increment$leftarrow $InitialCostValue/(2^{StepsToObstacle})$
@@
$MaxValue$leftarrow $max.~{}value~{}in~{}cells$ $C_{1}\dots C_{M}$
$MinValue$leftarrow $max.~{}value~{}in~{}cells$ $C_{1}\dots C_{M}$
for $z\leftarrow 1$ to $M$ do
@@
If we take the Figure 12 presented previously, if we consider the first step where we evaluate the costmap for the obstacles, this means:
@@
Figure 2: : Costmap calculation steps for each of the five considered obstacles (shown in red in each step), final value of costmap (in red) and normalized values (in blue)
The final costmap (the one in red) in Figure 2 shows a set of values that in some sense measure the impact of nearby obstacles in each cell. Normalization has been used to optimize the encoding of values to avoid wasting bits.
In a real environment, sensors have limitations and their reach is limited. For our analysis here, if we consider that the sensors of the robot can perceive their environment up to a distance of two cells (sensors can not perceive objects beyond two steps in the grid), using the Algorithm 1, we can represent what the robot would perceive in some position in the grid.
@@
Figure 3: : Costmap calculation steps for each detected obstacle (marked in red). In gray, obstacles which are beyond detection, hence they are not used in the calculation. Final value of costmap (in red) as the sum of all values for detected obstacles, and normalized values (in blue)
Now the localization problem is transformed into trying to match what the robot is perceiving. Such match is performed using with the previously stored costmap of the environment.
@@
Figure 4: : The robot localization requires to match the perception costmap (left), with the previously encoded and stored costmap (right)
@@
String searching algorithms, also known as string matching algorithms, are a crucial type of string algorithms designed to locate the position of one or more specific strings (referred to as patterns) within a larger string or block of text, which are commonly used in many different areas (Zhang, Zhaoyang (2022)): from daily activities (such as searching a text in an email), to language translation, data compression, or bioinformatics (when trying to find a certain DNA sequence (Sri et al. (2018))), etc.
In the case presented in the previous paragraphs (finding "323" in a longer text), a simple and straightforward (but naive) algorithm would be to create a loop and compare each element in the occupancy grid array with the target data. Starting with the first character of the larger string (the haystack), we compare it with the smaller string (the needle). If there is no match, we move to the second character and repeat the process, continuing this way for each position. Typically, only a few characters need to be checked at each position before determining if it is incorrect, so on average, this approach requires O(n + m) steps, where n is the length of the haystack and m is the length of the needle. However, in the worst-case scenario, such as when searching for a pattern like "aaaab" within "aaaaaaaaab", the process can take up to O(nm) steps.
Different algorithms have tried to optimize the time required for substring search in classical computing: from the optimized algorithms created in the seventies (such as the Knuth-Morris-Pratt Algorithm (Knuth et al. (1977)), or the Boyer-Moore Algorithm (Boyer and Moore (1977))), to the new approaches in the eighties (with the Yaeza-Bates (Baeza-Yates and Gonnet (1989)) and many others) and later. But this problem still requires O(n+ m) time if tackled classically (using classical computation).
As a complementary way of dealing with substring search, as any algorithmic improvements in this problem will result in great impacts in many areas, since 2000 quantum algorithms have been designed for the string matching problem (Ramesh and Vinay (2000)).
The starting point was the quantum Grover algorithm which searches for an element in an unordered database in O($\sqrt{n}$) time. As indicated by Hariharan and Vinay (2003), finding whether the pattern matches somewhere in the text is equivalent to searching in an unordered database; so we can have a quadratic speed-up compared to classical methods.
@@ 2.2 Using the quantum advantage in the localization problem
With the increasing demand for robust and accurate robot localization, researchers have been exploring innovative approaches to address the limitations of traditional methods. In recent years, the intersection of robotics and quantum computing has opened up new avenues for solving complex problems.
Inspired by the principles of quantum mechanics, we propose a novel approach to robot localization that leverages the power of quantum algorithms. Specifically, we draw inspiration from the Grover algorithm, a quantum search algorithm known for its exponential speedup over classical search algorithms.
We begin by discretizing the environment into a rectangular grid consisting of $nxn$ cells, each representing a small region of space. This grid serves as a framework for representing the robot’s surroundings and can be used to create an occupancy grid. In this occupancy grid, each cell is assigned a measurement derived from the robot’s onboard sensors (such as lidar, sonar, infrared readings…). The measurement represents the likelihood of the cell being occupied by an obstacle or being free space. This process is repeated as the robot explores the environment, resulting in a global annotated occupancy grid that captures the spatial layout of the environment.
When the robot is placed in an unknown position, it takes its new sensor readings and compares them to the annotated occupancy grid. To accelerate the search for the robot’s location, instead of using a classical method (such as using a particle filter) we employ a quantum search algorithm inspired by Grover’s algorithm. This algorithm efficiently searches the occupancy grid to find the cell that best matches the current sensor readings, effectively localizing the robot within the environment. By leveraging the power of quantum computing, the search process is significantly accelerated, allowing for fast and accurate robot localization even in complex and dynamic environments.
As an example, let’s use a simple environment divided into a a 4x4 grid as shown in the next Figure 5:
@@
Figure 5: : The black boxes mark the obstacles, and ®  marks the robot position

### v2

@@ 1 Introduction
Localization is a vital aspect of mobile robotics, enabling robots to navigate their environment efficiently and avoid obstacles. Without localization, mobile robots would not be able to determine their position and orientation, making it difficult to plan a path or make informed decisions about their movement ((Olson, 2000)). Localization allows mobile robots to create an internal map of their environment, which is essential for tasks such as surveying, manipulation, inspection, and delivery ((Huang and Lin, 2023)). In fact, based on the knowledge of a previously built map, localization is what enables mobile robots to perform tasks autonomously, allowing them to make the right actions and movements without human intervention.
@@
The quality of localization is heavily dependent on the generation of accurate maps, which is a computationally intensive task. Probabilistic localization methods, such as the Adaptive Monte Carlo localization (AMCL) algorithm, have been widely used in mobile robotics due to their accuracy and robustness ((Kristensen and Jensfelt, 2003)). However, these methods can be computationally demanding, especially when dealing with large maps or high-resolution sensor data. AMCL, in particular, uses a combination of sensor data and prior map knowledge to determine the probable location of a robot on a given map, but its computation complexity is proportional to the area of the grid of the map ((Alshikh Khalil and Hatem, 2022)).
@@
Recently, the integration of light detection and ranging (LiDAR) sensors has improved the accuracy of localization methods, but computational requirements remain a challenge ((Huang and Lin, 2023)). To overcome this challenge, researchers have started exploring the application of quantum computing in robotics, which has the potential to revolutionize the field by providing faster and more efficient computation ((Marek, 2020)).
@@
Quantum computing is a new paradigm for computing that has shown significant promise in various fields, including cryptography, optimization, and machine learning. Its application in robotics is still in its early stages, but it has the potential to provide breakthroughs in areas such as localization, planning, and control ((Petschnigg et al., 2019)). Quantum-inspired algorithms, such as Grover’s search, have been shown to provide computational speedups in certain tasks, making them an attractive solution for robotic applications ((Tandon et al., 2017)). Different proposals have been subject to research, such as planning ((Chella et al., 2023)), but new approaches that benefit from the quantum advantage have not yet been fully explored.
@@
This paper aims to evaluate the introduction of quantum computing in robotics, focusing on the application of quantum algorithms in robot localization. Specifically, it explores the use of Grover’s search algorithm in AMCL to improve the efficiency of localization in mobile robots. Although previous studies have explored the application of quantum planners in robotic navigation, their main focus has been on the use of quantum algorithms to improve decision-making optimization rather than integrating quantum computing into the localization module itself ((Chella et al., 2022)). Our approach aims to bridge this gap by introducing a pragmatic approach to robot localization that can be effectively implemented in real-world scenarios.
@@ 1.1 Research Objectives
Quantum computing has the potential to bring significant benefits to robotics science by enhancing computational power, optimizing algorithms, and enabling new types of processing that are not possible with classical computers ((Montanaro, 2015)). Quantum algorithms leverage the principles of quantum mechanics to perform certain computations more efficiently than classical algorithms.
Grover’s algorithm is one of the most famous quantum algorithms introduced by Lov Grover in 1996 ((Grover, 1996)). It is a quantum search algorithm that finds an element in an unsorted database of N items in $O(\sqrt{N})$ time, compared to classical algorithms that require $O(N)$ time. Grover’s algorithm is suitable for applications involving unsorted database search, cryptography, and combinatorial optimization, where it can provide a quadratic speedup over classical algorithms ((Guo, 2023)).
@@ 2 Background
Effective navigation is a fundamental capability for mobile robots, based on the successful integration of four essential components: sensor-based perception, self-localization, decision-making, and precise motion control. Among these components, self-localization is crucial, as it enables the robot to determine its position and orientation, providing a comprehensive understanding of its spatial disposition within its environment.
@@
Although classical-computing localization methods have shown evident and significant progress, the increasing complexity of robotic applications demands more efficient and robust techniques. The emergence of quantum computing offers possible new opportunities to enhance localization methods, with the promise of enabling faster processing and/or more accurate positioning. In the near future, quantum computing is expected to revolutionize the field of robotics in general (and localization in particular), enabling robots to navigate and interact with their environment.
@@ 2.1 Map encoding
Map encoding (encoding the map information into a form that can be manipulated by the robot’s CPU) is a required step in any localization technique. In recent decades, robotic research has explored and developed different mapping technologies ((Thrun, 2003)), ((Lluvia et al., 2021)), being the "occupancy grid" approach one of the dominant paradigms in the robotics field ((Collins et al., 2007)).
The core concept of the occupancy grid is to model the environment as a regularly spaced grid, where cells contain a certain numeric value that indicates the probability that the cell is traversable (measuring in certain way whether the cell is occupied by an obstacle, the cell has a nearby obstacle, or is fully traversable) (thrunoccupancy03).
In some naive approaches, the cost map utilized for computing the occupancy grid is binary (a cell cost is not zero when the cell is occupied by an obstacle and zero otherwise) (buhmann2000), but usually in cost map planning, the cost of each cell ranges from 0 to $\inf$, where 0 corresponds to a fully traversable cell, $\inf$ corresponds to an obstacle (untraversable), and any finite value in between corresponds to some degree of traversability (or some other cost function) (Brad96). In this paper, we follow the second approach, proposing a simple algorithm (simple enough so it can be used in a quantum computer) that creates a costmap that tries not only to identify the specific location of obstacles, but also the presence of nearby obstacles.
The proposed algorithm for the cost map has three steps: a first step where for every obstacle its influence in nearby cell is computed is presented, a second step where for each cell a sum of such influences is calculated (considering the distance to the obstacles), and a final third step where the cost values are normalized to compact the information (use less bits in its representation). It should be noted that distances are measured using the Manhattan distance (Suwanda20).
These steps are presented in Algorithm 1:
Algorithm 1 Simple costmap algorithm, assuming InitialCostValue=4
$B_{1}\dots B_{K}$ as a list of K obstacles
$C_{1}\dots C_{N}$ a a list of N cells in a $n\times n$ grid, being $N=n\times n$
@@
$Costmap~{}value~{}of~{}cell~{}C_{1}\dots C_{N}\leftarrow 0$
for $k\leftarrow 1$ to $K$ do
@@
         $Increment$leftarrow $InitialCostValue/(2^{StepsToObstacle-1})$
@@
$MaxValue$leftarrow $max.~{}value~{}in~{}cells$ $C_{1}\dots C_{N}$
$MinValue$leftarrow $max.~{}value~{}in~{}cells$ $C_{1}\dots C_{N}$
for $z\leftarrow 1$ to $N$ do
@@
Consider the setup shown in Figure 1.
Figure 1: Sample 4x4 grid with obstacles (in black)
If we consider the first step where we evaluate the costmap for the obstacles, this means:
@@
Figure 2: Costmap incremental calculation steps for each of the five considered obstacles (shown in red in each step), final value of costmap (in red) and normalized values (in blue)
The final costmap (the one in red) in Figure 2 shows a set of values that, in some sense, measure the impact of nearby obstacles in each cell. Normalization has been used to optimize the encoding of values to avoid wasting bits.
In a real environment, sensors have limitations, and their reach is limited. For our analysis here, if we consider that the robot sensors can perceive their environment up to a distance of two cells (sensors cannot perceive objects beyond two steps in the grid), using the Algorithm 1, we can represent what the robot would perceive in some position on the grid.
@@
Figure 3: The robot is into an unknown location. Figure shows costmap calculation steps for each detected obstacle (marked in red). In gray, obstacles which are beyond detection, hence they are not used in the calculation. Final value of costmap (in red) as the sum of all values for detected obstacles, and normalized values (in blue)
Now, the localization problem transforms into trying to match what the robot is perceiving. This match is performed using the previously stored costmap of the environment.
@@
Figure 4: The robot localization requires to match the perception costmap (left), with the previously encoded and stored costmap (right)
@@
String searching algorithms, also known as string matching algorithms, are a crucial type of string algorithms designed to locate the position of one or more specific strings (referred to as patterns) within a larger string or block of text, which are commonly used in many different areas ((Zhang, Zhaoyang, 2022)): from daily activities (such as searching for a text in an email), to language translation, data compression, or bioinformatics (when trying to find a certain DNA sequence ((Sri et al., 2018))), etc.
In the case presented in the previous paragraphs (finding "323" in a longer text), a simple and straightforward (but naive) algorithm would be to create a loop and compare each element in the occupancy grid array with the target data. Starting with the first character of the larger string (the haystack), we compare it with the smaller string (the needle). If there is no match, we move to the second character and repeat the process, continuing this way for each position. Typically, only a few characters must be checked at each position before determining whether it is incorrect, so on average, this approach requires $O(n+m)$ steps, where n is the length of the haystack and m is the length of the needle. However, in the worst-case scenario, such as when searching for a pattern like "aaaab" within "aaaaaaaaab", the process can take up to O(nm) steps.
Different algorithms have tried to optimize the time required for substring search in classical computing: from the optimized algorithms created in the seventies (such as the Knuth-Morris-Pratt Algorithm ((Knuth et al., 1977)), or the Boyer-Moore Algorithm ((Boyer and Moore, 1977))), to the new approaches in the eighties (with Yaeza-Bates ((Baeza-Yates and Gonnet, 1989)) and many others) and later. But this problem still requires O(n+ m) time if tackled classically (using classical computation).
@@ 2.2 Using the quantum advantage in the localization problem
As a complementary way of dealing with substring search, as any algorithmic improvements in this problem will result in great impacts in many areas, since 2000 quantum algorithms have been designed for the string matching problem ((Ramesh and Vinay, 2000)).
The starting point was the quantum Grover algorithm that searches for an element in a unordered database in $O(\sqrt{n})$ time. As indicated by (Hariharan and Vinay, 2003), finding whether the pattern matches somewhere in the text is equivalent to searching in an unordered database; so we can have a quadratic speed-up compared to classical methods.
With the increasing demand for robust and accurate robot localization, researchers have explored innovative approaches to address the limitations of traditional methods. In recent years, the intersection of robotics and quantum computing has opened new avenues for solving complex problems.
Inspired by the principles of quantum mechanics, we propose a novel approach to robot localization that takes advantage of the power of quantum algorithms. Specifically, we draw inspiration from the Grover algorithm, a quantum search algorithm known for its exponential speedup over classical search algorithms.
We begin by discretizing the environment into a rectangular grid consisting of $n\times n$ cells, each representing a small region of space. This grid serves as a framework for representing the robot’s surroundings and can be used to create an occupancy grid. In this occupancy grid, each cell is assigned a measurement derived from the robot’s onboard sensors (such as lidar, sonar, infrared readings, etc.). The measurement represents the likelihood that the cell is occupied by an obstacle or is free space. This process is repeated as the robot explores the environment, resulting in a global annotated occupancy grid that captures the spatial layout of the environment.
When the robot is placed in an unknown position, it takes its new sensor readings and compares them with the annotated occupancy grid. To accelerate the search for the robot’s location, instead of using a classical method (such as using a particle filter), we employ a quantum search algorithm inspired by Grover’s algorithm. This algorithm efficiently searches the occupancy grid to find the cell that best matches the current sensor readings, effectively localizing the robot within the environment. Taking advantage of quantum computing, the search process is significantly accelerated, allowing for fast and accurate robot localization even in complex and dynamic environments.
As an example, let’s use a simple environment divided into a 4x4 grid as shown in the Figure 5:
@@
Figure 5: The black boxes mark the obstacles, and ®  marks the robot position